OpenStudy (anonymous):

Explain in complete sentences how you would use the graphing method to solve the following system of equations. Provide the solution to the system and explain what the solution represent in the graph. 3x+6y=18 2x+5y=20 HELP

OpenStudy (anonymous):

you graph the two equations on a plane, and where both lines intersect, that should be your answer (x,y).

OpenStudy (anonymous):

3x+6y=18 -- -- -- 3 3 3 x+2x=6

OpenStudy (anonymous):

i mean x+2y=6

OpenStudy (anonymous):

x+2y-6=0

OpenStudy (anonymous):

he has to use the graphing method though. but that is one way of solving it... making them equal each other :)

OpenStudy (anonymous):

oh okay. what formula are you talking about?